Sunday, October 11, 2015

Lab hours have been extended

Starting Monday, October 12th, the Open Computing Facility lab's hours have been extended.

They are now:

  • Monday: 9am-6pm, 7pm-9pm
  • Tuesday-Thursday: 9am-9pm
  • Friday: 9am-8pm
  • Saturday/Sunday: 11am-7pm
See the lab page for a full list of hours and holidays.

Many thanks to the Publications Center staff for extending the lab's hours!

Friday, October 09, 2015

Scheduled outage Tuesday 10/13 from 7am to 7:30am

There will be about 30 minutes of network downtime on Tuesday, October 13th from 7am to 7:30am while IST performs firewall maintenance. All OCF services will be unavailable during this time.

Sunday, September 20, 2015

Service outage due to hanging server reboot (resolved)

Public-facing services are down. Web hosting and SSH access have been down since approximately 2am PDT. The login server (tsunami) was restarted at 10:21pm.

Staff are investigating the outage.

Update 3:13am: Services are now recovering and should be fully restored within a few minutes.

Update 3:18am: All services should now be restored. Please contact if you are still encountering trouble.

Saturday, September 19, 2015

Lab closed, some services affected by campus-wide network downtime

There was a fire in the campus datacenter a few hours ago which took down CalNet and related campus websites. See for more details.

The following were affected by the CalNet outage, but are no longer affected

  • Account creation
  • Password resets

WiFi access in the lab is still down. All OCF services themselves are fine.

The Publications Center made the decision to close the lab due to the WiFi outage. It will reopen once WiFi has been restored.

Sunday, August 30, 2015

Removing automatic redirects from www.ocf.b.e/user -> www.ocf.b.e/~user

We are removing the automatic redirects from www.ocf.b.e/user -> www.ocf.b.e/~user.

There appears to be virtually no traffic to these, so you should not notice any changes. will continue to work.

This will take effect in one week (Sept. 6).

Saturday, August 08, 2015

Downtime Wednesday for security updates

All servers will be restarted the night of Wednesday, August 12th to apply security updates. Sorry for the inconvenience.

Update: The downtime took longer than expected when our file server failed to come back up within the ~5 minutes we expect for a restart. It took an additional ~30 minutes to run a fsck. We'll be more careful to watch for and schedule these at more convenient times in the future.

Saturday, June 20, 2015

Ongoing downtime due to server crash

At 1:07pm today, hal, our primary production server, froze. We are on-site and working to restore it. We are moving the important servers to another machine while we investigate, as hal continues to experience issues.

Update 2:07pm: Service has been restored, but we are continuing to move servers to a different machine. There will be some downtime as we continue the migration, but it will affect single services only.

Remaining to migrate: (updated 5:04pm)

  • firestorm (ldap)
  • death (www)
  • pestilence (dns, dhcp)
  • supernova (admin)
  • maelstrom (mysql)
  • tsunami (ssh)
  • anthrax (smtp)
  • sandstorm (group smtp)
  • biohazard (apphost)
  • lightning (puppet)
  • earthquake (accounts)
  • typhoon (rt)
  • blight (wiki)
  • flood (irc)
  • reaper (jenkins)
  • dev-earthquake (dev-accounts)
  • pollution (cups)

Update 5:20pm: All VMs are migrated to jaws, and all services should be restored. We'll be debugging and rebuilding hal in the near future, and will be scheduling downtime some time in the next few weeks to move VMs back. We'll post a followup here when we have a date in mind.